What is the height,correct to the nearest cm,of a 5 litre cylindral oil container with the radius of 20cm?
Answer by Cromlix(4381)

5 litres = 5000cm^3
Volume of a cylinder = Pi*r^2*h
5000 cm^3 = Pi*20^2*h
Height = 5000/Pi*20^2
Height = 3.97 cm
Height = 4 cm.
Hope this helps.
:-)